LPC meeting summary 18-07-2016 - final

Minutes overview      LPC home

Minutes and Summary

Main purpose of the meeting: Discussion of the remaining 2.5km β* commissioning.

Introduction (Christoph Schwick)

Christoph briefly summarised the last week of LHC operation. The availability of the machine continued to be outstanding. The machine has been in Stable Beams for 70% of the time during the last week. So far 16/fb have been delivered to ATLAS and CMS in 71 days of pp data taking. There are still 80 days of pp data taking foreseen in the schedule.

Christoph summarised the first experience with BCMS beams. Three small problems have been encountered and solved so far:

A first step of decreasing the controlled emittance blow up in the PS has been performed leading to lower emittances in the LHC and hence higher luminosity. Beams remained stable.

Christoph reported that on Wednesday the LMC will decide on the possibility to shift the Abort Gap Keeper (AGK) to a later position in the orbit, so that more space for injecting bunch trains will be available. This shift is possible since this year the longest injection is limited to 144b due to the vacuum problem in the SPS. This will lead to an increase of 7% colliding bunches in ATLAS/CMS and around 15% in LHCb/ALICE. Christoph estimated that the total luminosity for 2016 would increase by about 3% provided that we continue to collect 1.5/fb to 1.75/fb per week during the remaining pp physics time.

Finally Christoph asked the experiments when they would be able to discuss how they cope with the new beam conditions in their physics analysis. A meeting will be scheduled once the experiments will be able to comment on this.

Totem's view for the 2.5km β* commissioning and requirements for the physics run (Jan Kaspar)

Jan reminded that Totems main goal for the 2.5km β* run is the precision measurement of the cross section in the Coulomb-Nuclear interference region, to be possibly able to discriminate between different models. He estimated that at least 100µb-1 would be needed to achieve this goal. Witold Kozanecki asked if this estimate takes into account that the luminosity at β* = 2.5km is about 2.5 times less than in the previous run with β* = 1km. Jan answered that this factor was not included in the estimate.

During the physics run Totem would like to have one non colliding bunch pair which could have a smaller intensity than the colliding pairs. In order to optimise for high luminosity Totem prefers that the remaining intensity budget (a total of 3.0 x 1011 protons per beam) is distributed onto two colliding bunch pairs. Totem prefers low emittances to keep the beam divergence low.

Jan stated that the optics from March 2016 is good for Totem. During the summary he underlined once more the importance of high statistics and hence a possibly long data taking period.

During the discussion Helmut Burkhard emphasised that this run is performed under very particular conditions and that it is very hard to predict how this run will turn out. He reminded that during the 1km β* run in 2012 one of the two beams suffered from blow up, and that the beam had to be scraped down twice during the data taking, reducing its intensity by a factor of 1/2.

It was confirmed that the run will be taken in ADJUST mode and Stable Beams will not be declared.

After the meeting the LPC discussed with the MPP group in order to find out that it is NOT possible to go beyond the limit of 3x1011 protons for the total beam intensity. There are hard-coded limits in the machine protection system which would not allow to mask certain interlocks when this intensity is superseded. However, the interlock to dump the beam when a bunch is passing the threshold of the BPMs and is becoming invisible to the protection system during an ongoing fill may be masked in these special conditions. This could be considered if the non colliding bunch will be chosen significantly less intense than the colliding bunches.

Alfa's view for the 2.5km β* commissioning (Patrick Fassnacht)

After having thanked the Machine experts for the successful commissioning of the standard 2.5km optics Patrick stated that the optics as it is today is disfavoured by Alfa since one of their vertical pots is positioned at a location where the phase advance is almost precisely 180 degree which would make it impossible to use the so called "subtraction method" to determine the kinematic variable t in the collisions. However this method is crucial for the physics analysis of the data.

Alfa requests a change of the optics implemented as an add-on to the existing and commissioned optics which changes the phase advanced slightly at the vertical pot in question. A solution for this modification exists on paper. Two fills would be required to implement the change.

Alfa estimates the number of hours to implement this change with 8-12 hours (2 fills). This would be needed in addition to the time required for background studies.

Since the improved optics would only change the phase advance at the pots by ≈1 degree, it was questioned to which precision the phase advanced can be know/measured. The conclusion of the discussion was that in the past the phase advance was exactly where it was predicted and therefore Alfa is having confidence that the current optics is not suitable for their physics programme, but that the alternative solution mentioned would solve their problem.

It was clarified that such an optics modification would be local and would not change the optics for Totem.

On a question from Alexander Oh, Patrick stated that the shortcoming of the current optics could not be compensated with statistics (i.e. by taking more data).

In the discussion Helmut Burkhard underlined that the run is pushing many parameters to the limit and that it is a priori not obvious that the modification will succeed. It has to be tried out and if found to be not possible, one would have to revert to the already commissioned optics.

During the discussion on when to schedule this optics development it was asked of the background measurements could be done before the optics change or if they would have to be done/repeated with the final optics. No final conclusion to answer this question was found, however Alfa estimates that the background measurements if done before the optics change, would remain valid.

It was concluded that, in order to provide conditions for Alfa which allow for best possible physics results, the optics change should be tried, however the scheduling is complicated due to availability of experts, and remains to be worked out. The LPC regretted that this requirement for the optics was brought to their attention at such a late stage in the preparation of this run.

Greg Rakness asked if the schedule of the MD1 would change in case the 2.5km development session would be postponed. Experiments would like to know about any change as soon as possible to allow for efficient planning of the MD time. An answer to this question could not be given yet.